# If It’s Not Enough, Make It So: Reducing Authentic Data Demand in Face Recognition through Synthetic Faces

### v1

@@ IV-D Augmentation Impact on Combined Training Data (RQ4)
Our results show that applying data augmentation on both authentic and synthetic subsets (RAug-All settings) led to overall lower average accuracy, in comparison to the cases in which no data augmentation was applied. On average, models trained after augmenting all the identities part of the training dataset, when WF constituted the authentic component, were impacted by $-0.01\%$ ($+0.54\%$), $+0.23\%$ ($-0.43\%$), $-0.11\%$ ($-0.43\%$), $+0.21\%$ ($+0.10\%$), $-0.18\%$ ($-0.92\%$) and $-0.14\%$ ($-0.22\%$) w.r.t. the baselines trained on not augmented data. Each comparison was made by taking into consideration combined and authentic training sets involving 1K, 2K, 4K, 6K, 8K and 10K authentic identities, respectively. A slight but consistent improvement w.r.t. the combined baselines is instead observable after augmenting only the synthetic training subset (RAug-Synt settings). In this scenario, FR models trained on authentic identities subsets from WF combined with augmented synthetic samples from DC (GC) led to $-0.01\%$ ($+0.31\%$), $+0.49\%$ ($-0.30\%$), $+0.10\%$ ($+0.03\%$), $+0.27\%$ ($+0.52\%$), $-0.06\%$ ($0\%$) and $-0.06\%$ ($0.17\%$) changes in average verification accuracy.
@@
Given the consistent enhancements observed by augmenting only the synthetic samples in our combined datasets, it became interesting to explore whether this behavior is confirmed across different sources of authentic identities. To address this, we extended our investigation to training FR models on synthetic data combined with authentic identities from M2-S. Results are reported in Table VI and VII. In this scenario, it can be observed that combining augmented GC data with authentic identities from M2-S constantly led to worse verification performances than the combined baselines, which were impacted by $+0.23\%$, $-2.49\%$, $-4.83\%$, $-4.09\%$, $-3.38\%$ and $-4.42\%$ while involving 1K, 2K, 4K, 6K, 8K and 10K authentic identities, respectively. Combining augmented DC data with the same authentic identities instead, always led to well noticeable improvements for every authentic data subset. Specifically, FR models trained with augmented DC samples achieved $1.32\%$, $1.00\%$, $1.11\%$, $0.94\%$, $0.68\%$ and $0.93\%$ improvements. The consistence of our results while training FR models on augmented DC samples further reinforces our previous considerations, pointing to end-to-end synthetic data augmentation as a promising direction for performance improvement in FR when using combined data.
@@ V Conclusions and Future Work
To conclude, we connect the insights coming from individual analysis and discuss the results and future research.
Our initial analysis aimed to comprehend the performance differences between FR models exclusively trained on authentic data and those trained solely on synthetic data generated through SOTA generative methods. This investigation led to a better understanding of the impact of data sources variation on model performance, under an equal number of identities across datasets. Our findings align consistently with prior research [2, 27], confirming that FR models trained on the a large, full authentic dataset exhibit the highest performance, achieving an average accuracy of $94.63\%$ across the benchmarks in our experiments.
@@
In our third analysis, our goal was to assess whether a training dataset with a fixed number of synthetic identities and a progressively increasing number of authentic ones could attain accuracy levels, at the very least, comparable to a larger dataset comprising solely authentic identities. For both authentic datasets, all settings involving combined datasets yielded superior outcomes compared to the synthetic-only baselines, showcasing improvements of approximately $30\%$ in several cases. Moreover, training FR models on synthetic data combined with subsets of authentic identities consistently led to improvements compared to models trained exclusively on authentic subsets. As another observation, models trained on combined datasets that included synthetic data from DC consistently outperformed those involving synthetic data from GC, especially in cross-pose benchmarks.

### v4

@@ IV-D Impact of data augmentation
The reported results demonstrated that augmenting both authentic and synthetic subsets (RAug-All settings) used to train FR models leads to overall lower average verification accuracy w.r.t. the case in which no augmentation is applied to the training data. On average, the verification accuracy of models trained on a fully augmented training dataset (with subsets from WF as authentic identities), was impacted by $-0.01\%$, $+0.23\%$, $-0.11\%$, $+0.21\%$, $-0.18\%$ and $-0.14\%$ while including synthetic identities from DC and by $+0.54\%$, $-0.43\%$, $-0.43\%$, $+0.10\%$, $-0.92\%$ and $-0.22\%$ while including synthetic identities from GC. Each comparison was made by taking into consideration combined and authentic training sets involving 1K, 2K, 4K, 6K, 8K, and 10K authentic identities, respectively.
A slight but consistent improvement, w.r.t. the baselines trained on data with no augmentation, is achieved by augmenting only the synthetic training subset (RAug-Synt settings). In this scenario, FR models trained on authentic identity subsets from WF combined with augmented synthetic samples from DC led to $-0.01\%$, $+0.49\%$, $+0.10\%$, $+0.27\%$, $-0.06\%$ and $-0.06\%$ changes in average verification accuracy. Similarly, training FR models on authentic identity subsets from WF combined with augmented synthetic samples from GC led to $+0.31\%$, $-0.30\%$, $+0.03\%$, $+0.52\%$, $0\%$, $+0.17\%$ improvements in term of average verification accuracy.
@@
Given the consistent enhancements observed by augmenting only the synthetic samples in our combined training datasets, we investigated whether this behavior is confirmed across different sources of authentic identities. To address this, we extended our investigation to training FR models on synthetic data combined with authentic identities from M2-S. Results are reported in Table VI and VII. In these settings, it can be observed that FR models trained on augmented GC data combined with authentic identities from M2-S constantly led to slightly lower verification performances than the baselines, which were impacted by $+0.23\%$, $-2.49\%$, $-4.83\%$, $-4.09\%$, $-3.38\%$ and $-4.42\%$ while involving 1K, 2K, 4K, 6K, 8K and 10K authentic identities, respectively. FR models trained with augmented DC samples achieved $1.32\%$, $1.00\%$, $1.11\%$, $0.94\%$, $0.68\%$ and $0.93\%$ improvements. The consistency of our results while training FR models on augmented DC samples combined with authentic identities further reinforces our previous considerations. This points out that augmenting synthetic data is a promising direction for improving the FR performances, in case the FR model is trained on combining subsets of synthetic and authentic data.
@@ V Conclusions and Future Work
To conclude, we connect the insights coming from individual analysis and discuss the results and future research. Our initial analysis aimed to comprehend the performance differences between FR models exclusively trained on authentic data and those trained solely on synthetic data generated through SOTA generative methods. This investigation led to a better understanding of the impact of data source variation on model performance, under an equal number of identities across datasets. Our findings align consistently with prior research [44, 28], confirming that FR models trained on the large, full authentic dataset exhibit the highest performance, achieving an average accuracy of $94.63\%$ across the benchmarks in our experiments.
@@
In our third analysis, we assessed whether a training dataset with a fixed number of synthetic identities and a progressively increasing number of authentic ones could attain accuracy levels, at the very least, comparable to a larger dataset comprising solely authentic identities. For both authentic datasets, all settings involving combined datasets yielded superior outcomes compared to the synthetic-only baselines, showcasing improvements of approximately $30\%$ in several cases. Moreover, training FR models on synthetic data combined with subsets of authentic identities consistently led to improvements compared to models trained exclusively on authentic subsets. As another observation, models trained on combined datasets that included synthetic data from DC consistently outperformed those involving synthetic data from GC, especially in cross-pose benchmarks.